# Human Skin Permittivity Characterization for Mobile Handset Evaluation at Sub-THz

### v1

@@ Abstract
This manuscript proposes a method for characterizing the complex permittivity of the human finger skin based on an open-ended waveguide covered with a thin dielectric sheet at sub-terahertz frequencies. The measurement system is initially analyzed through full-wave simulations with a detailed finger model. Next, the model is simplified by replacing the finger with an infinite sheet of human skin to calculate the forward electromagnetic problem related to the permittivity characterization. Following this, a radial basis network is employed to train the inverse problem solver. Finally, the complex permittivities of finger skins are characterized for 10 volunteers. The variations in complex relative permittivity across different individuals and skin regions are analyzed at 140 GHz, revealing a maximum deviation of ±$\pm 0.7$ for both the real and imaginary parts. Repeated measurements at the same location on the finger demonstrate good repeatability with a relative estimation uncertainty <±$<\pm 1\%$.
@@ III-B S-parameter Measurements
The measurement setup is illustrated in Fig. 4. During our measurements, the fingers under consideration were not washed or dried to maintain a typical daily humidity level encountered during mobile terminal usage. In the context of finger-antenna interactions, particularly during mobile terminal usage, it is the fingerprint sides that notably influence the radiation patterns of the mobile terminal antennas. Consequently, the measurements primarily focus on the fingerprint sides, while the fingernail sides are not considered.
@@
Additionally, different parts of the fingers, especially the fingerprint regions, may display varying permittivities. Hence, averaging permittivity estimates from multiple finger measurements is essential for obtaining representative values. To collect finger permittivity estimates across various individuals, we conducted measurements with 10 volunteer subjects aged between 18 and 40 years old. Each volunteer underwent 10 measurements, focusing on the fingertips of their index, middle, and ring fingers.
@@ IV Results and Discussions
The mean of the estimated relative permittivities is depicted in Fig. 5 based on 10 measurements for each volunteer. For instance, at $140~{}\rm GHz$, the real part of the relative permittivity ranges from $4.0$ to $5.3$, while the imaginary part varies from $1.1$ to $2.4$. In most cases, the real parts fall within the range of $4.3$ to $4.8$, while the imaginary parts fall between $1.4$ to $2.1$. There are slight variations in the permittivity values among the subjects, with a maximum deviation of ±$\pm 0.7$ for both the real and imaginary parts. The average relative permittivity of all volunteers is $\varepsilon_{\rm r}=4.6-1.7\rm i$ at $140~{}\rm GHz$.
Figure 5: Estimated real and imaginary parts of the finger relative permittivity for 10 subjects. Mean values based on 10 measurements are presented.
Table 1: Estimation uncertainty and permittivity variation of tissue over $10$ measurement samples at $140~{}\rm GHz$
Real part Imaginary part
Estimation Uncertainty ±$\pm 0.32~{}\%$$\pm 0.95~{}\%$
Volunteer 1 $\left[3.80,4.30\right]$ $\left[1.05,1.70\right]$
Volunteer 2 $\left[4.29,5.16\right]$ $\left[0.86,1.40\right]$
Volunteer 3 $\left[3.75,4.25\right]$ $\left[1.78,2.55\right]$
Volunteer 4 $\left[4.56,5.30\right]$ $\left[1.88,2.59\right]$
Volunteer 5 $\left[4.96,5.72\right]$ $\left[1.91,2.90\right]$
Volunteer 6 $\left[4.84,5.70\right]$ $\left[1.02,1.96\right]$
Volunteer 7 $\left[4.07,4.70\right]$ $\left[1.56,2.10\right]$
Volunteer 8 $\left[4.24,4.50\right]$ $\left[1.30,1.69\right]$
Volunteer 9 $\left[4.01,5.10\right]$ $\left[1.15,1.90\right]$
Volunteer 10 $\left[4.17,4.85\right]$ $\left[1.77,2.51\right]$
To assess estimation uncertainty, 10 repeated measurements were conducted within a short duration of 2 minutes at the same finger location. The estimated relative deviation of permittivity is within ±$\pm 0.32\%$ of the mean value for the real parts and ±$\pm 0.95\%$ for imaginary parts, as presented in Table 1. These relative deviations are smaller than those reported in [gao2018towards], indicating the high repeatability of the measurement method. Additionally, a variation in permittivity for each volunteer was observed repeating the measurements 10 times for the same finger. For the majority of volunteers, measuring the same finger 10 times resulted in variations of <±$<\pm 0.4$ for both real and imaginary parts.
@@ V Conclusions
In this manuscript, we propose a novel method for characterizing human skin permittivity using an open-ended waveguide at sub-THz frequencies. The necessary contact area of the human skin under test on the sheet is identified through full-wave simulations so that the antenna-hand interaction model can excellently emulate practical measurements. The full-wave simulations provide reflection coefficients to train the RBN, which is applied to characterize finger permittivities of living humans. The estimated relative permittivity shows smaller uncertainty than ±$\pm 1\%$ for both real and imaginary parts, demonstrating the high repeatability of the proposed approach. The measured complex relative permittivity estimates are valuable base for studying finger-antenna interactions of sub-THz mobile handsets. Furthermore, the proposed method can be applied to characterize the permittivity of various parts of human skin.

### v2

@@ Abstract
This manuscript proposes a method for characterizing the complex permittivity of the human finger skin based on an open-ended waveguide covered with a thin dielectric sheet at sub-terahertz frequencies. The measurement system is initially analyzed through full-wave simulations with a detailed finger model. Next, the model is simplified by replacing the finger with an infinite sheet of human skin to calculate the forward electromagnetic problem related to the permittivity characterization. Following this, a radial basis network is employed to train the inverse problem solver. Finally, the complex permittivities of finger skins are characterized for 10 volunteers. The variations in complex relative permittivity across different individuals and skin regions are analyzed, revealing a deviation of <±$<\pm 1.5$ for both the dielectric constants and loss factors across 140 to 220 GHz. Repeated measurements at the same location on the finger demonstrate good repeatability with a relative estimation uncertainty <±$<\pm 1.5\%$.
@@ III-B S-parameter Measurements
The measurement setup is illustrated in Fig. 4.
@@
During our measurements, the fingers under consideration were not washed or dried to maintain a typical daily humidity level encountered during mobile terminal usage. In the context of finger-antenna interactions, particularly during mobile terminal usage, it is the fingerprint sides that notably influence the radiation patterns of the mobile terminal antennas. Consequently, the measurements primarily focus on the fingerprint sides, while the fingernail sides are not considered.
Additionally, different parts of the fingers, especially the fingerprint regions, may display varying permittivities. Hence, averaging permittivity estimates from multiple finger measurements is essential for obtaining representative values. To collect finger permittivity estimates across various individuals, we conducted measurements with 10 volunteer subjects aged between 18 and 40 years old, and their heights are from 1.60 to 1.85 meters. Each volunteer underwent 10 measurements, focusing on the fingertips of their index, middle, and ring fingers.
@@ IV Results and Discussions
The mean of the estimated relative permittivities is depicted in Fig. 5a for each volunteer. For instance, at $140~{}\rm GHz$, the dielectric constant (real part) ranges from $4.0$ to $5.9$, while the loss factor (imaginary part) varies from $1.9$ to $3.9$. In most cases, the dielectric constants fall within the range of $3.9$ to $4.8$, while the loss factors fall between $1.9$ to $2.4$. The average relative permittivity of all volunteers is $\varepsilon_{\rm r}=4.7-2.4\rm i$ at $140~{}\rm GHz$. It can be seen that Volunteer 7 and 8 show higher dielectric constants and loss factors and show different trends with frequency changes. This is because these two volunteers’ hands are more moist.
Figure 5: Mean dielectric constant and loss factor based on 10 different finger parts for a) each volunteer and b) all the volunteers.
Figure 6: Measurement uncertainty of dielectric constant and loss factor based on 20 measurements at the same finger location for a) Volunteer 1 and b) Volunteer 2.
Figure 7: (a)-(j) Dielectric constant and loss factor based on 10 different finger parts for 10 volunteers.
To assess measurement uncertainty, 20 repeated measurements were conducted within a short duration of 2 minutes at the same finger location of two volunteers. The estimated relative deviation of permittivity is within ±$\pm 1.4\%$ of the mean value for the dielectric constants and within ±$\pm 1.2\%$ for loss factors, as presented in Fig. 6. These relative deviations are smaller than those reported in [gao2018towards], indicating the high repeatability of the measurement method. Additionally, a variation in permittivity for each volunteer was observed based on the 10 measurements on different finger parts shown in Fig. 7. The majority of volunteers’ fingers exhibit a variation of <±$<\pm 0.5$ for both the dielectric constants and loss factors. Some volunteers’ dielectric constants show a larger variation, like Fig. 7d7e, and 7g, while some of the others show a larger variation in loss factors, like Fig. 7b7c, and 7h.
@@ V Conclusions
In this manuscript, we propose a novel method for characterizing human skin permittivity using an open-ended waveguide at sub-THz frequencies. The necessary contact area of the human skin under test on the sheet is identified through full-wave simulations so that the antenna-hand interaction model can excellently emulate practical measurements. The full-wave simulations provide reflection coefficients to train the RBN, which is applied to characterize finger permittivities of living humans. The estimated relative permittivity shows smaller uncertainty than ±$\pm 1.5\%$ for both the dielectric constants and loss factors, demonstrating the high repeatability of the proposed approach. The measured complex relative permittivity estimates are valuable base for studying finger-antenna interactions of sub-THz mobile handsets. Furthermore, the proposed method can be applied to characterize the permittivity of various parts of human skin.